# Ozawa’s Intersubjectivity Theorem as justification of RQM’s postulate on internally consistent descriptions

### 11 Jan 2024 v1

@@ Abstract
The Ozawa’s Intersubjectivity Theorem (OIT) proved within quantum measurement theory supports the new postulate of relational quantum mechanics (RQM), the postulate on internally consistent descriptions. But from OIT viewpoint postulate’s formulation should be completed by the assumption of probability reproducibility.
@@ 1 Introduction
Recently Ozawa’s Intersubjectivity Theorem (OIT) [1] proved within quantum measurement theory (the indirect measurement scheme for quantum instruments theory[2, 3]) was used as objection to QBism [4], to its basic interpretational statement about the individual agent perspective on measurement’s outcomes [5]-[7]. This privacy component of QBism was also criticized in my previous articles [8, 9], mainly from the probabilistic perspective. Now QIT provides the strong argument that measurement’s outcome is intersubejctive and, although one can still interpret it as a private experience, such interpretation loses its elegance. QBists replied to my objection with two papers [10, 11] (the second one presents more elaborated counterargument). The detailed response two these papers will be prepared by Ozawa.
The discussions on the intersubjectivity problem in QBism stimulated my interest to this problem for other information interpretations of QM. A reviewer of one of my papers pointed that relational quantum mechanics (RQM) [12] was suffering of the same problem. It seems that it cannot be resolved within “old RQM”. Its resolution required invention of an additional postulate on internally consistent descriptions (postulate 6 in [13]). By treating this postulate within quantum instrument theory, one immediately see that OIT supports this postulate. So, in contrast to QBism which was challenged by OIT, RQM’s postulate 6 is mathematically justified via OIT.
@@
We start with a brief recollection of the basics of RQM and the problem of intersubjectivity in its framework. Here we follow article [13]. Since I am not an expert in RQM, I shall mainly cite this article.
@@ Acknowledgments
This study was partially supported by the EU-project: CA21169 - Information, Coding, and Biological Function: the Dynamics of Life (DYNALIFE).

### 13 Mar 2024 v2

# Relational Quantum Mechanics: Ozawa’s Intersubjectivity Theorem as justification of the postulate on internally consistent descriptions

@@ Abstract
The Ozawa’s Intersubjectivity Theorem (OIT) proved within quantum measurement theory supports the new postulate of relational quantum mechanics (RQM), the postulate on internally consistent descriptions. But from OIT viewpoint postulate’s formulation should be completed by the assumption of probability reproducibility. We remark that this postulate was proposed only recently to resolve the problem of intersubjectivity of information in RQM. In contrast to RQM for which OIT is a supporting theoretical statement, QBism is challenged by OIT.
keywords: Relational Quantum Mechanics, Ozawa Intersubjectivity Theorem, postulate on internally consistent descriptions, quantum instruments, measurement process, reproducibility of probability distribution, unitary evolution
@@ 1 Introduction
Recently Ozawa’s Intersubjectivity Theorem (OIT) [1] proved within quantum measurement theory (e.g., [2]-[8]) was used as objection to QBism [9], to its basic interpretational statement about the individual agent perspective on measurement’s outcomes [10]-[12]. (This privacy component of QBism was also criticized in my previous articles [13, 14], mainly from the probabilistic perspective.) Now QIT provides the strong argument that measurement’s outcome is intersubejctive and, although one can still interpret it as a private experience, such interpretation loses its elegance. QBists replied to my objection with two papers [15, 16] (the second one presents more elaborated counterargument).
In a few words the QBism’s counterargument is that compatible measurement devices $M_{1}$ and $M_{2}$ considered in OIT should be related not to two distinct observers $Q_{1}$ and $O_{2}$ as was done in my article [9], but to one observer $O$ who performs these two measurements. As was clearly explained in papers [15, 16], this interpretation of such measurements is consistent with the general ideology of QBism. The important impact of article [9] to QBism is that this article attracted the attention (at least some) of QBists to quantum measurement theory [2]-[8]. QBism intensively explores one of the basic elements of this theory - POVMs. But QBists didn’t employ the important part of the quantum measurement theory, namely, that every measurement described by completely positive instrument has a parallel process described by an interaction between the object and the apparatus, or more mathematically every abstract measuring process determines its completely positive instruments and that every completely positive instruments arises in this way. (This was proven by Ozawa in article [3].)
The discussions on the intersubjectivity problem in QBism stimulated my interest to this problem for other information interpretations of QM. A reviewer of one of my papers pointed that relational quantum mechanics (RQM) [17]- [21] was suffering of the same problem. It seems that it cannot be resolved within “old RQM”. Its resolution required invention of an additional postulate on internally consistent descriptions (postulate 6 in [21]). By treating this postulate within quantum instrument theory, one immediately see that OIT supports this postulate. So, in contrast to QBism which was challenged by OIT, RQM’s postulate 6 is mathematically justified via OIT.
@@
As well as QBism, RQM can earn (for the moment, it is not clear how much?) from its framing withing the mathematical formalism of quantum measurement theory [2]-[8]. Besides the present paper, an attempt to proceed in this direction was done in article [22]. However, the conclusions of this article are debatable. I am not an expert in RQM and I am not able to evaluate whether the authors of [22] correctly understood the conceptual premises of RQM.
We start with a brief recollection of the basics of RQM and the problem of intersubjectivity in its framework. Here we follow article [21].
@@ Acknowledgments
This study was partially supported by the EU-project: CA21169 - Information, Coding, and Biological Function: the Dynamics of Life (DYNALIFE). I would like to thank Masanao Ozawa for numerous lessons on quantum measurement theory, discussions on intersubjectivity of quantum measurements, and hospitality during my visits to Nagoya during the last 20 years.